1. |x | = 6

2. |3x + 2| = 14

3. - 5| x + 1| = -10

4. |x - 2| + 10 = 12

5. |x| = - 5

6. |2x + 6| - 4 = 20

7. 6 - 3|2x + 6| = 0

8. 10 - |x + 2| = 12
: 1. |x | = 6

2. |3x + 2| = 14

3. - 5| x + 1| = -10

4. |x - 2| + 10 = 12

5. |x| = - 5

6. |2x + 6| - 4 = 20

7. 6 - 3|2x + 6| = 0

8. 10 - |x + 2| = 12

Answer by Nate(3495)

There are two answers: solving normally, and solving when you negate
1. | x | = 6
x = 6 and x = -6
2. |3x + 2| = 14
3x + 2 = 14 and 3x + 2 = -14
3x = 12 and 3x = -16
x = 4 and x = -16/3
3. -5| x + 1 | = -10
| x + 1 | = 2
x + 1 = 2 and x + 1 = -2
x = 1 and x = -3
... and so on ...